Note Receivable: On Jan 1st, 2014, Xena Inc. provided services in exchange for a 2-year $100,000, 8% note receivable that pays interest quarterly on March 31st, June 30th, September 30th, and Dec 31st. The customer’s normal borrowing rate (market rate) is 12%.

On, Jan 1st, 2014, the carrying value of the note receivable is $___________ Round to the nearest dollar.

a. Prepare a well-labeled schedule (with debits/credits shown) for the journal entries through the life of the Note.

b. Prepare the original Journal Entry to record the issue of the Note Receivable

c. Prepare the Journal Entry to record the Interest on 12/31/15 for the Note Receivable
Answer #1

А в C F G H I $ 100,000 $ 2,000 face value of note quarterly interest payments total quarterly periods quarterly market interB19 date account titles and explanation debit credit Jan 1st, 2014 $ 100,000 note receivable discout on note receivable servi

for formulas and calculations, refer to the image below -

D face value of note quarterly interest payments total quarterly periods quarterly market interest rate 100000 =D2*8%/4 =2*4B19 date account titles and explanation debit credit Jan 1st, 2014 note receivable discout on note receivable service revenue

In case you have any query, kindly ask in comments.